OpenStudy (anonymous):

INTEGRATE : x^3/(x^2+1)^2

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ x ^{3} }{ (x ^{2 }+1)^{2} }\]

OpenStudy (kainui):

I would probably use this substitution: \[x=\tan \theta\]

OpenStudy (anonymous):

ok, let me try it

OpenStudy (solomonzelman):

Substitute u for x^2 ?

OpenStudy (anonymous):

uhmmm, i thought i was going to easily figure the tan theta, but i am stuck

OpenStudy (kainui):

Actually you could probably do x^2+1=u and that will be MUCH easier.

OpenStudy (anonymous):

ok, let me try that here

OpenStudy (anonymous):

\[u= x ^{2} + 1\]

OpenStudy (anonymous):

\[du =2xdx\] \[\frac{ 1 }{ 2 }du=xdx\] from above \[x ^{2}=u-1\] \[\int\limits_{}^{}\frac{ u-1 }{ u ^{2} }\]

OpenStudy (anonymous):

is this the correct pathway

OpenStudy (kainui):

You got it. =) I am only cringing a little bit when you don't include your dx's and du's on your integrals but overall you're totally right!

OpenStudy (anonymous):

thank you very much, i will surely remember to add them!

OpenStudy (anonymous):